OpenStudy (anonymous):

Find the x-intercepts for the parabola defined by this equation: y=3x^2+3x-6 Write your answer as two ordered pairs: (x1,y1), (x2,y2)

OpenStudy (anonymous):

the x-intercepts are when Y equals zero, so you simply set Y to zero, then solve the equation for when it equals zero. Here's the process: 0=3x^2+3x-6 0=3(x^2+x-2) 0=x^2+x-2 0=(x+2)(x-1) Note: yes, this is a bit of a jump here, but you just have to factor the polynomial, and you'll get this answer. x=-2 or x=1 (-2,0) (1,0)

OpenStudy (anonymous):

3x^2 + 3x - 6 = 0 3(x^2 + x - 2) = 0 3(x - 1)(x + 2) = 0